Fermi National Laboratory
 

Operation Summary for May 12, 2010 - May 14, 2010

 

Wednesday May 12

The day shift began with Operations monitoring the Tevatron (TeV) store 7807 (172.11E30) and the Recycler stash (130.7E10), with the Antiproton Source (Pbar) stacking (15.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI in access, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations resumed sending beam to NuMI at 2:03 PM. A Tevatron separator sparked and caused a small loss in luminosity. A Linac RF station low level RF reboot caused a small amount of downtime.

During the evening shift, Operations terminated store 7807 at 7:35 PM. Operations put the TeV into HEP shot setup and established store 7808 at 8:58 PM with an average initial luminosity of 360.0E30. Operations established beam to the Meson MTest beam line briefly before it was disabled at 6 PM. System personnel investigated multiple Booster RF trips.

Thursday May 13

The midnight shift began with Operations monitoring store 7808 (190.77E30) and the stash (120.8E10), with Pbar stacking (7.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that store 7808 aborted at 6:05 AM when a lightning strike caused the TeV hardwire loop interlock to drop out; Operators put the TeV into recovery mode. Shortly after 4 AM, Operations reported multiple lightning strikes that caused several minor problems. Pbar was hit the hardest. It lost a beam momentum cooling system that required consultation with system personnel. Operators prepared the Pbar rings for access.

During the day shift, the Tevatron was in recovery and the Pbar Accumulator was in access. Operations put the TeV into shot setup at 9:41 AM. Operations established store 7811 at 10:44 AM with an AIL of 293.3E30. Operations disabled the Booster beam at 10:07 AM to prepare for a Booster and MI access to repair damage caused by the lightning strikes. Pbar came out of access at 12:24 PM. The Booster and MI accesses ended in the late afternoon. Operations established beam to MiniBooNE by at 3:09 PM. Operations established beam to NuMI at 3:51 PM.

During the evening shift, Operations reported that they had to delay beam due to two Pbar power supplies that required local resets. Operations established beam to Pbar at 4:47 PM. Operations established beam to MTest at 4:46 PM and disabled it at 7:42 PM. Controls problems caused a few minutes of NuMI downtime.

Friday May 14

The midnight shift began with Operations monitoring store 7811 (64.0E30) and the stash (162E10), with Pbar stacking (13.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations struggled with Booster RF sparking at the start of shift; operators reported finding a problem with a RF station modulator power supply. Operators turned the station off and tuned the RF. Operators brought the Booster intensity back to normal at 2:30 AM. Operations reported several minor Linac, Booster, and Main Injector RF trips.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/14/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory